how many electrons could have the following set of quantum numbers

(a) n = 3. l = 2 

_______ electrons 

(b) n=4, l=2, ml=1 

_______ electrons
Answer #1

a)

For given value of n and l,

Number of orbitals = 2l+1

= 2*2+1

= 5

Each orbital can have 2 electrons.

So, number of electrons = 5*2 = 10

Answer: 10

b)

n, l and ml together can represent only 1 orbital.

It can have 2 electrons.

Answer: 2